OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6).

OpenStudy (anonymous):

\[Y = \frac{ 1 }{ 3 }x + 7\] \[Y = -3x + 3\] \[Y = \frac{ 1 }{ 3 }x + 3\] \[Y = -3x + 7\]

OpenStudy (anonymous):

Because the original slope is -3 and we are looking for a parallel slope the slope for the second equation will be -3. y - 6 = -3(x - (-1)) y - 6 = -3x + 1 +6 = + 6 y = -3x 7

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = -3x + 7

OpenStudy (anonymous):

It's wrong. O.o

OpenStudy (anonymous):

Hold on. Gonna check it.

OpenStudy (anonymous):

Ya, it's wrong. y = -3x + 3. Sorry. Did you submit it yet?

OpenStudy (anonymous):

Yeah, but its good. :)

OpenStudy (anonymous):

Oh, okay. Sorry again.

OpenStudy (anonymous):

It's alright!